let a, b be the roots of the equation x^2-10cx-11d=0 and c,d be the roots of x^2 -10ax-11b=0, Find the value of a+b+c+d.

let  a, b be the roots of the equation  x^2-10cx-11d=0 and c,d be the roots of x^2 -10ax-11b=0, Find the value of a+b+c+d.

Grade:11

5 Answers

Rahul askIITiansExpert.IITR
13 years ago
As a + b = 10c and c + d = 10a
ab = -11d , cd = -11b
ac = 121 and (b + d) = 9(a + c)
a2 - 10ac - 11d = 0
c2 - 10ac - 11b = 0
a2 + c2 - 20ac - 11(b + d) = 0
(a + c)2 - 22(121) - 11 × 9(a + c) = 0
(a + c) = 121 or -22 (rejected)
a + b + c + d = 1210.

Difficult to understand - post it here and we will get you the answer and detailed solution very quickly. We are all IITians and here to help you in your IIT JEE preparation.

DO APPROVE IF THE ANSWER HELPS YOU.

Now you can win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Regards,

Askiitians Experts

Rahul- IIT Roorkee
Kushagra Mishra
8 years ago
how is ac=121?
I dont get it xD :/
Premanshu Rao
8 years ago
Why did we reject -22?
Akshay
8 years ago
you will not reject -22, a=-11,b=-99,c=-11,d=-99 is also a possible solution. Just substitute and check
Rishabh kumar
5 years ago
a b c d must be distinct numbers so that (-22) is rejected in the real question of iit jee it was written that a ,b,c,d must be distinct